Given the mass of iron nucleus as 55.85u and A=56, find the nuclear density?

Grade:11

1 Answers

suraj sarma
6 years ago
According to rutherford`s relationsradius of nucleus = ro×A13 ro = constant = 1.2×10−15 mthereforer = 1.2×10−15 ×56−−√3r = 4.59 ×10−15 malso,1u = 1.67×10−27 kg55.85 u = 9.327 ×10−26 kg = mass of nucleus of iron atomnuclear density = mass of nucleus of iron atomvolume of nucleus nuclear density = 9.327 ×10−26 kg43π×r3 =9.327 ×10−26 kg43π×(4.59 ×10−15 )3nuclear density = 4.851×1018 kg/m^3